An investment will pay $150 at the end of each of the next 3 years, $250 at the end of Year 4, $300 at the end of Year 5, and $500 at the end of Year 6. If other investments of equal risk earn 11% annually, what is its present value? Its future value?
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

=150/1.11+150/1.11^2+150/1.11^3+250/1.11^4+300/1.11^5+500/1.11^6

=$976.6(Approx)

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period

Future value=976.6*(1.11)^6

=$1826.64(Approx).